### v1

@@ Appendix B Other Generalization Frameworks
Rademacher complexity (Bartlett & Mendelson, 2002) exactly measures the ability for a model to fit uniform $\{+1,-1\}$ random noise. In particular, the Rademacher complexity of a hypothesis space $\mathcal{H}$ and an input sample $\{x_{i},\dots,x_{n}\}$ is $\mathcal{R}(\mathcal{H})=\mathbb{E}_{\sigma}\left[\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}h(x_{i})\right]$, where $\sigma_{i}$ are i.i.d. Rademacher random variables ($\{+1,-1\}$ with equal probability). The expected risk of a hypothesis $h$ is then bounded as $R(h)\leq\hat{R}(h)+2\mathcal{R}(\mathcal{H})+C$, where $C$ is a constant defined by the loss function, $n$, and the confidence $1-\delta$ of the bound. Thus, if the model has a hypothesis space $\mathcal{H}$ that can fit the Rademacher noise, then the Rademacher generalization bound will be uninformative.
@@
Rademacher $\mathcal{R}_{n}(\mathcal{H})$ $R(h)\leq R^{S}(h)+2\mathcal{R}_{n}(\mathcal{H})+3\sqrt{\frac{\log(2/\delta)}{2n}}$ 2000s Measures expected maximum correlation any $h\in\mathcal{H}$ can achieve with uniform $\{+1,-1\}$ samples. Does not explain overparametrization, benign overfitting, or double descent.
@@ Appendix C Countable Hypothesis Bound
$R(h)\leq\hat{R}(h)+\Delta\sqrt{\frac{\log 1/P(h)+\log 1/\delta}{2m}}.$ (5)
@@
As $m\hat{R}(h)$ is the sum of independent and bounded random variables, we can apply Hoeffding’s inequality (Hoeffding, 1994) for a given choice of $h$. For any $t>0$
@@ Appendix D Experimental Details
Figure 1(f) is adapted from Maddox et al. (2020) and uses a ResNet-18 with increasing layer width, measures train loss, test loss, and effective dimensionality for $\alpha=1$. For Figure 1(g) we use the random feature least squares model $Xw=y$ with each column of $X_{i}=y_{i}+\epsilon$ where $\epsilon\sim\mathcal{N}(0,1)$. We measure MSE, and use $\alpha=1$ for effective dimensionality.

### v2

@@ Appendix B Other Generalization Frameworks
Rademacher complexity (Bartlett & Mendelson, 2002) exactly measures the ability for a model to fit uniform $\{+1,-1\}$ random noise. In particular, the Rademacher complexity of a hypothesis space $\mathcal{H}$ and an input sample $\{x_{i},\dots,x_{n}\}$ is $\mathcal{R}(\mathcal{H})=\mathbb{E}_{\sigma}\left[\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}h(x_{i})\right]$, where $\sigma_{i}$ are i.i.d. Rademacher random variables ($\{+1,-1\}$ with equal probability). The expected risk of a hypothesis $h$ is then bounded as $R(h)\leq\hat{R}(h)+2\mathcal{R}(\mathcal{H})+C$, where $C$ is a constant defined by the loss function, $n$, and the confidence $1-\delta$ of the bound. Thus, if the model has a hypothesis space $\mathcal{H}$ that can fit the Rademacher noise, then the Rademacher generalization bound will be uninformative — unless it is adapted to include a prior that assigns higher density to certain solutions over others, much like how we move from a standard finite hypothesis bound with a uniform prior over hypotheses to a countable hypothesis bound with arbitrary prior in Appendix C.
@@
Rademacher $\mathcal{R}_{n}(\mathcal{H})$ $R(h)\leq\hat{R}(h)+2\mathcal{R}_{n}(\mathcal{H})+3\sqrt{\frac{\log(2/\delta)}{2n}}$ 2000s Measures expected maximum correlation any $h\in\mathcal{H}$ can achieve with uniform $\{+1,-1\}$ samples. Does not explain overparametrization, benign overfitting, or double descent.
@@ Appendix C Countable Hypothesis Bound
$R(h)\leq\hat{R}(h)+\Delta\sqrt{\frac{\log 1/P(h)+\log 1/\delta}{2n}}.$ (5)
@@
As $n\hat{R}(h)$ is the sum of independent and bounded random variables, we can apply Hoeffding’s inequality (Hoeffding, 1994) for a given choice of $h$. For any $t>0$
@@ Appendix D Experimental Details
Figure 1(f) is adapted from Maddox et al. (2020) and uses a ResNet-18 with increasing layer width, measures train loss, test loss, and effective dimensionality for $\alpha=1$. Similar to Maddox et al. (2020), in Figure 1(g) we use the random feature least squares model $Xw=y$ with each column of $X_{i}=y_{i}+\epsilon$ where $\epsilon\sim\mathcal{N}(0,1)$. We measure MSE, and use $\alpha=10$ to compute the effective dimensionality of the parameter covariance matrix (inverse Hessian).